### v1

# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

Abstract
Despite significant advancements in text generation and reasoning, Large Language Models (LLMs) still face challenges in accurately performing complex arithmetic operations. To achieve accurate calculations, language model systems often enable LLMs to generate code for arithmetic operations. However, this approach compromises speed and security and, if finetuning is involved, risks the language model losing prior capabilities. We propose a framework that enables exact arithmetic in a single autoregressive step, providing faster, more secure, and more interpretable LLM systems with arithmetic capabilities. We use the hidden states of an LLM to control a symbolic architecture which performs arithmetic. Our implementation using Llama 3 8B Instruct with OccamNet as a symbolic model (OccamLlama) achieves 100% accuracy on single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), outperforming GPT 4o and on par with GPT 4o using a code interpreter. OccamLlama also outperforms both Llama 3 8B Instruct and GPT 3.5 Turbo on multistep reasoning problems involving challenging arithmetic, thus enabling small LLMs to match the arithmetic performance of even much larger models. We will make our code public shortly.
1 Introduction
Since the release of GPT 3, Large Language Models (LLMs) have dramatically improved in their text generation and reasoning capabilities. This has enabled success in downstream applications including machine translation [1, 2], sentiment analysis [3, 4, 5], and interactive dialogue generation [6], with language models even surpassing human experts on some academic benchmarks that require reading comprehension, reasoning and coding [7]. However even industry-leading LLMs such as GPT 4 cannot reach 100% accuracy on simple arithmetic [8], limiting their ability to perform basic mathematical tasks. This hinders potential applications of LLMs ranging from chat-bot physics tutors to LLM-powered automated research which would accelerate scientific discovery and technological innovation. The poor arithmetic performance of LLMs is particularly acute for small LLM agents, limiting their usage in smartphone or in multi-agent applications.
To enable accurate calculations, language model systems often resort to running code written by an LLM. However, this comes at the cost of speed; the model must perform multiple autoregressive steps to generate code that performs the appropriate arithmetic operations. This increased decoding time may negatively impact applications such as multi-agent workflows [9, 10] where speed is essential. At the same time, code-based LLM arithmetic mechanisms may increase system vulnerability by providing a mechanism for arbitrary LLM-generated code execution.
We propose an alternative, a framework which enables exact and interpretable LLM arithmetic in a single autoregressive step, providing faster and more secure arithmetic capabilities in LLM systems. Our framework uses the hidden states of an LLM to control a symbolic architecture that performs arithmetic. Although our method can in principle work with any symbolic architecture, in this paper we use an interpretable neurosymbolic architecture known as OccamNet [11, 12] because of its interpretability and scalability. Because of this, we term our method OccamLLM, or OccamLlama when using a Llama model as the LLM.
Our core contributions are as follows:
  1. 1.
    We develop a framework for exact and interpretable LLM arithmetic in a single autoregressive step without catastrophic forgetting [13] or vulnerability from code generation. We explore how to train OccamLlama, including data generation, decoder architecture, and loss function.
  2. 2.
    We benchmark OccamLlama on arithmetic tasks, demonstrating that OccamLlama achieves 100% accuracy on arbitrary single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), more than double the accuracy of GPT 4o. OccamLlama performs slightly better than GPT 4o with Code Interpreter while using on average more than 50x fewer tokens to generate.
  3. 3.
    We benchmark on mathematical reasoning problems, showing that OccamLlama can sustain long generations. OccamLlama outperforms both Llama 3 8B Instruct and GPT 3.5 Turbo on reasoning problems with challenging arithmetic.
Figure 1: The OccamLLM system. For each autoregressive step, the language model hidden states for that token are fed into a decoder block which assigns weights to OccamNet. The system feeds the most recent numbers from the text into OccamNet, which then evaluates the sparse function specified by its weights. The decoder then determines whether to use the LLM output or the OccamNet output.
2 Related Work
Arithmetic Performance in LLMs.
Prior research has trained models on synthetic data, finding that such models can achieve near-perfect accuracy on addition [14, 15], subtraction [15], multiplication [14, 15], division [15], and raising to powers [15]. Furthermore, these prior models have been tested only on arithmetic datasets, so their generality has not been assessed.
Other work focuses on finetuning LLMs which are already trained on large amounts of general-purpose data on math datasets. Both full-parameter [16, 17] and parameter-efficient (PEFT) [18] finetuning strategies have been applied in this regard. However, finetuning on a single dataset carries the risk of catastrophic forgetting of an LLM’s previously acquired linguistic skills [19]. While PEFT techniques have been shown to partially mitigate this effect, this area is still one of active research [20, 21].
LLMs with Tool Use.
Another thrust of prior research has focused on LLM tool use, which we believe is most directly related to our methods. Calc-X [22] introduces a technique to offload arithmetic computations to an external tool like a calculator. The authors curated a large dataset of arithmetic problems and trained a language model that learns to interact with a calculator through the use of tags to signify the calling of the external tool. Several other works [23, 24, 25] follow a similar idea, using crowd workers to annotate tool calls and using this data to train language models to interact with external tools such as a web searching tool, a calculator, or a translation system. These approaches can be prohibitively expensive in annotation costs; Toolformer [26] overcomes this cost by using in-context learning and a language model to generate datasets containing the necessary ‘API’ tool calls via a self-supervised loss. Further, the above methods all require finetuning of the LLM, placing the LLM at risk of losing generality and its original language modelling abilities through catastrophic forgetting. In contrast, our approach does not involve training the language model. Our ‘external tool’ is a symbolic neural network which can be trained to correctly use the hidden states of the language model to perform the required arithmetic computations. The language model is kept frozen throughout this process. Unlike other tool-calling approaches, where the cost of data annotation to train for tool-calling interaction can be prohibitively expensive, in our method, each task only requires manually annotating tens of prompts, a high annotation efficiency. Other prior methods leverage prompt engineering to improve arithmetic performance of LLMs; this is done either through chain-of-thought [27], or to encourage LLMs to use a code interpreter [28, 29, 30]. Contrary to these methods, our approach does not use code kernels; this provides several advantages: 1) it enables tool use without expending compute on autoregressive steps for token generation, and 2) it avoids running potentially incorrect or malicious code generated by language models.
Table 1: OccamLLM is the only approach to improving the arithmetic capabilities of a pretrained LLM which 1) enables single pass arithmetic, 2) does not risk catastrophic forgetting from finetuning, 3) does not require arbitrary code execution, and 4) provides an interpretable process.
No Catastrophic No Arbitrary
Single Pass Forgetting Code Execution Interpretable
Fine Tuning
Tool Use
OccamLLM
3 Methods
3.1 OccamLLM: Combining a Language Model with a Symbolic Model
In short, the OccamLLM system combines a language model with a symbolic model, namely OccamNet, that can perform arithmetic operations like addition and subtraction. For each token, the corresponding internal hidden states of the language model are fed into a decoder module which initializes the symbolic model so that it executes the operation required by the task described in the input text. A string parser feeds the necessary numbers from the text into OccamNet, which evaluates the desired expression. Finally, a decoder determines whether to use the language model output or the OccamNet output for generating the next token.
In the example shown in Figure 1, a decoder determines how to initialize OccamNet from the language model hidden states, choosing to have OccamNet perform addition. The text parser then feeds the numbers $6$ and $7$ into the symbolic model, which adds the numbers, returning $13$ Finally, a decoder decides to use the OccamNet output instead of the language model output, so the system outputs 13. The new sentence, including the 13, is tokenized and fed back to the LLM to continue normal autoregressive generation. The language model might later generate “Since she ate two apples, she now has,” at which point the switch will again trigger OccamNet, this time implementing $13-2$ and returning $11$
In the subsections below we describe the OccamLLM system which we find performs substantially better than other LLMs. For a much larger analysis of alternate architectures and losses, see Appendix B.
3.1.1 OccamNet
OccamNet is a symbolic architecture that provides an interpretable way of parametrizing probability distributions over a space of functions [11]. We leave a more thorough explanation of OccamNet to [11], describing only the relevant components here. For a more detailed explanation, see Appendix C.
An $l$-layer OccamNet with primitives $\mathcal{P}$ and $n$ inputs is an architecture that defines a probability distribution over the space of functions representable as compositions of the primitives in $\mathcal{P}$ up to depth $l.$ For example, a two-layer OccamNet with primitives $\mathcal{P}=\{\sin,\cos\}$ and one input represents a probability distribution over the set
=$\mathcal{F}=\{x,\sin(x),\cos(x),\sin(\sin(x)),\sin(\cos(x)),\cos(\sin(x)),\sin(\sin(x))\}.$
OccamNet has the structure of an $n$-input, $l$-internal-activation-layer multilayer perceptron with the biases removed and the activations in each layer replaced by the primitives $\mathcal{P},$ as shown in Figure 2a. Activation functions may have multiple inputs. We rename the linear layers softmax layers, denote the weights of the $i$th softmax layer as $\mathbf{W}_{i}$, and denote the combined weights of OccamNet as $\mathbf{W}.$
We define the probability distribution which OccamNet parametrizes by specifying how to sample from it. For each softmax layer output node (shown in red in Figure 2), we select a single connection to that node from a softmax layer input node by sampling from the distribution given by the softmax of the weights of the connections to the different inputs. This process produces a directed acyclic graph (DAG) defining a computational path through the OccamNet activations, such as the one shown in Figure 2b. In this way, each DAG represents a function on the inputs of OccamNet.
Figure 2: a) A schematic of the OccamNet architecture, with softmax layers in grey and their outputs in red. b) A Directed Acyclic Graph (DAG) (with edges not connected to the output removed for clarity) formed by sampling from OccamNet. This DAG corresponds to the function $\sin(\sin(x_{1})\cdot\exp(x_{0})).$ Modified from [11].
To ensure that OccamNet can represent all possible compositions of functions in $\mathcal{P}$ up to depth $l,$ we include the following modifications to the OccamNet architecture: 1) for each softmax layer, we concatenate its inputs with the previous softmax layer’s inputs to enable the representation of functions with fewer than $l$ compositions, and 2) we repeat primitives in the $i$th activation layer $A^{l-i}$ times, where $A$ is the maximum number of inputs of any of the primitives, to ensure that a sufficient number of each primitive is available at each layer. We refer to this modified architecture as complete OccamNet as it can represent the complete set of desired functions. The resulting architecture is shown in Figure 6 in the appendix.
In principle, OccamLLM can work with any symbolic model, i.e., any model that can parameterize a set of symbolic functions or a distribution over such functions. We choose OccamNet as opposed to, for example, a transformer [31] or recurrent neural network [32], for two reasons: 1) OccamNet is interpretable, which we hypothesize makes controlling OccamNet easy for a decoder to learn, and 2) OccamNet is parallelizable over multiple samples, allowing for scalable training.
3.1.2 OccamLLM Decoder
The OccamLLM decoder takes the hidden states of a language model and outputs an initialization for OccamNet. This gives the LLM control over which function to apply on the inputs. The decoder acts on each input token separately, producing a different OccamNet initialization for each. Therefore, the arithmetic operations predicted may change along an input sequence, allowing OccamNet’s use for different computations in a single example. This is crucial in multi-step reasoning scenarios where OccamNet is employed several times for different purposes.
Many decoder architectures are possible. We choose to parameterize the weights of each softmax layer of OccamNet independently, as $(\mathbf{W}_{1},\ldots,\mathbf{W}_{l})=(\text{Decoder}_{1}(\mathbf{h}),\ldots,\text{Decoder}_{l}(\mathbf{h})),$ where $\mathbf{h}$ are the hidden states of the language model. We choose
$\text{Decoder}_{i}(\mathbf{h})=\text{MLP}_{i}\left(\sum_{j=1}^{L}w_{i,j}\mathbf{h}_{j}\right)+\mathbf{W}_{i}^{(0)}$ (1)
where $\mathbf{h}_{j}$ are the hidden states of the $j$th layer of the language model, $w_{i,j}$ are trainable weights, $\text{MLP}_{i}$ are two-layer multilayer perceptrons (MLPs), and $\mathbf{W}_{i}^{(0)}$ are untrained weights which initialize all functions to have approximately equal probabilities according to the initialization described in [11] and explained in Appendix C.4.
3.1.3 OccamLLM Switch
We similarly train a decoder for a switch that, for each input token, is fed the hidden states of the language model and selects whether to use the output of OccamNet or the output of the language model. The decoder outputs a single number from 0 to 1, where all numbers less than or equal to 0.5 correspond to using the output of the language model and all numbers greater than 0.5 correspond to using the output of OccamNet. We choose the following architecture for the switch decoder:
$\text{Decoder}_{\text{switch}}(\mathbf{h})=\text{sigmoid}\left(\text{MLP}_{\text{switch}}\left(\sum_{j=1}^{L}w_{\text{switch},j}\mathbf{h}_{j}\right)\right).$ (2)
3.2 Data Generation
We create synthetic datasets to train the OccamLLM decoder, which contain instruction prompts for diverse arithmetic tasks. To generate datasets of arbitrary size, we create prompts with placeholders for numbers to sample from. Each prompt includes a string with number placeholders, the sampling value range for each number, and a function that computes the answer to the query given the sampled input numbers. The prompts fall into two main categories: purely arithmetic tasks and reasoning problems.
Purely arithmetic prompts are formed by expressions including only symbols, without any natural language added, such as “$3+85=$.” We create prompts using the following operations: +(,)+($\cdot$,$\cdot$)+ ( ⋅ , ⋅ ), (,)-($\cdot$,$\cdot$)- ( ⋅ , ⋅ ), ×(,$\times$($\cdot$,$\cdot$)× ( ⋅ , ⋅ ), ÷(,$\div$($\cdot$,$\cdot$)÷ ( ⋅ , ⋅ ), $\text{sqrt}(\cdot)$, $\text{power}(\cdot,\cdot)$, $\log_{e}(\cdot)$, $\exp(\cdot)$, $\sin(\cdot)$, and $\cos(\cdot)$.
We also include word problems that require one or two reasoning steps. We generated 150 single step word problems and 40 multi-step reasoning problems which we modified from examples in the MultiArith training dataset [33].
3.2.1 OccamNet Decoder Training Data
For training the decoder that controls the weights of OccamNet, we created two types of examples, single queries and concatenated queries. For single queries, we select a single prompt from the problems generated as discussed in Section 3.2. We use the Llama 3 8B Instruct chat template and fill in the query as the user input and the result as the assistant response, prepending “Answer = ” to the later in randomly selected samples (see Appendix A.1 for further details). For the concatenated queries of examples, we select a random number of prompts and concatenate the query-response pairs without using the Llama 3 8B Instruct chat template. The OccamNet decoder is trained to predict only the results of the last query in the sequence. This strategy helps OccamLLM to learn which operation to perform without becoming confused by the text that appears before, which is useful for continuous generation. To create the training dataset, each example is sampled by first randomly selecting the kind of example, then randomly selecting the type(s) of prompt(s) used, and finally randomly sampling the input values from the range corresponding to each selected prompt.
3.2.2 OccamLLM Switch Training Data
To train the switch, we generate examples of possible LLM outputs for given input expressions, and label the outputs with sequences of 0s or 1s corresponding to whether the language model or the OccamNet output should be used for the next token. Some input-output pairs correspond to the prompts described in Section 3.2. For this, the LLM output is set to “The answer is” or “Answer = ” and the label sequence only contains a 1 at the end to indicate the system should use OccamNet to compute the answer. We also manually created and labeled several other examples for diverse scenarios, to explicitly teach the system in which cases it should or should not use OccamNet. To create the dataset, we concatenate a random number of these user input - assistant output pairs in a conversational fashion, using the Llama 3 8B Instruct chat template.
3.3 OccamLLM Training
We train the OccamLLM decoder and the switch separately, as they do not share weights. In all cases, the weights of the LLM are kept frozen. In the first step, we train the system to predict the answer to examples generated by the method explained in Section 3.2.1. The OccamNet decoder needs to process the set of hidden states corresponding to the last token of the response and set the weights of OccamNet appropriately, so that the correct arithmetic expression is sampled. In this step, we use a rescaled REINFORCE [34] loss:
$\mathcal{L}(x,y;W)=-\frac{\sum_{f\sim p_{W}}R(f(x),y)\log p_{W}[f]}{\sum_{f\sim p_{W}}R(f(x),y)},$ (3)
where $p_{W}[f]\equiv\text{ON}(f;\text{Decoder}_{W}(\mathbf{h}(x)))$ is the probability distribution represented by the decoder-initialized OccamNet. We find setting $R(f(x),y)=1$ if $f(x)=y$, and 0 otherwise, most effective.
Minimizing this loss steers the decoder towards assigning higher probabilities to the functions that maximize the reward $R(f(x),y)$, which measures the similarity between the correct answer $y$ and the prediction of OccamNet $f(x)$. We discuss the OccamNet loss in more detail in Appendix B.
The second step involves training the decoder to route the outputs to OccamNet when needed. We train the switch decoder alone, freezing the weights of the OccamNet decoder of the previous step and minimizing the binary cross-entropy loss between the switch output and the desired output for each token. The OccamLLM switch decoder learns when to route the output to OccamNet in diverse contexts.
4 Experiments
For all OccamLLM results, we use Llama 3 8B Instruct [35] as the underlying language model. As such, we call our model OccamLlama. We use a 1 layer Complete OccamNet with primitives
$\mathcal{P}=\{+(\cdot,\cdot),-(\cdot,\cdot),\times(\cdot,\cdot),\div(\cdot,\cdot),\text{sqrt}(\cdot),\text{power}(\cdot,\cdot),\log_{e}(\cdot),\exp(\cdot),\sin(\cdot),\cos(\cdot)\}.$
This single layer OccamNet can be invoked by the LLM several times during generation to perform any arithmetic operation accurately. To use the trained OccamLlama for inference, we sample the highest probability function from OccamNet as described in Appendix C.3.
We benchmark our methods against unmodified Llama 2 7B Chat (Llama 2 7B) [36], unmodified Llama 3 8B Instruct (Llama 3 8B) [35], gpt-3.5-turbo-0125 (GPT 3.5 Turbo), gpt-4o-2024-05-13 (GPT 4o) [37], and gpt-4o-2024-05-13 with Code Interpreter (GPT 4o + Code). To reduce costs, for GPT 4o with Code Interpreter, we test a random subset of 200 datapoints for each dataset.
To determine if a model output is correct, we parse all numbers in the model output and if one of them “matches” the correct answer, we determine that the result is correct. We mark each correct result as 100% accuracy and each incorrect result as 0% accuracy. For each model on each dataset, we report the mean accuracy and the standard error of the mean. To determine if a number matches the result, we first determine how many places after the decimal $d$ the number should be accurate to. If the number is an integer, we set $d$ to 2. Otherwise, we set $d$ to the number of places after the decimal in the model output, clipped between 2 and 5. Finally we state that a number “matches” the result if the number and the result differ by less than $10^{-d}$. We present further experiment details, including hyperparameters, prompts, etc. in Appendix A.
4.1 Simple Arithmetic Problems
Table 2: Accuracy on arithmetic tasks. OccamLlama uses Llama 3 8B Instruct as the underlying language model. Higher is better. Bold indicates best performance for each row.
OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
7B Chat 8B Instruct Turbo Code
Addition 100.0$\pm$0.0 19.2$\pm$1.2 44.9$\pm$1.6 65.2$\pm$1.5 95.7$\pm$0.6 100.0$\pm$0.0
Subtraction 100.0$\pm$0.0 8.7$\pm$0.9 34.4$\pm$1.5 59.8$\pm$1.6 85.6$\pm$1.1 99.5$\pm$0.5
Multiplication 100.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 99.0$\pm$0.7
Division 100.0$\pm$0.0 2.8$\pm$0.5 35.3$\pm$1.5 10.7$\pm$1.0 38.6$\pm$1.5 100.0$\pm$0.0
Square Root 100.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.9$\pm$0.3 18.6$\pm$1.2 100.0$\pm$0.0
Exponential 100.0$\pm$0.0 0.3$\pm$0.2 3.1$\pm$0.5 12.5$\pm$1.0 23.2$\pm$1.3 100.0$\pm$0.0
Logarithm 100.0$\pm$0.0 0.1$\pm$0.1 0.0$\pm$0.0 17.1$\pm$1.2 21.3$\pm$1.3 100.0$\pm$0.0
Sine 100.0$\pm$0.0 7.6$\pm$0.8 7.0$\pm$0.8 13.4$\pm$1.1 39.3$\pm$1.5 100.0$\pm$0.0
Cosine 100.0$\pm$0.0 0.8$\pm$0.3 1.5$\pm$0.4 6.7$\pm$0.8 32.8$\pm$1.5 100.0$\pm$0.0
Average 100.0$\pm$0.0 4.4$\pm$0.2 14.0$\pm$0.4 20.7$\pm$0.4 39.5$\pm$0.5 99.8$\pm$0.1
[…]
The results of these evaluations are shown in Table 2. More detailed results, including relative error and results for 3 and 5 digit arithmetic, are shown in Appendix A. OccamLlama has $100.0\pm 0.0$% accuracy on all tasks, missing 0 out of 9000 problems. On the other hand, GPT 4o with Code Interpreter missed 3 out of the 1800 problems it faced (we tested on less problems to save cost) and had accuracy $99.8\pm 0.1$%.
Furthermore, GPT 4o with Code Interpreter generates on average more than 54 tokens to answer these problems, whereas our model uses OccamNet on the first forward pass. This means that, barring advanced decoding techniques such as speculative decoding [38], GPT 4o would need to be more than 50x faster than OccamLlama per forward pass to be comparable in overall generation speed on these tasks.
Figure 3: Accuracy of OccamLlama and baselines on reasoning tasks. Higher is better. OccamLlama achieves better performance than Llama 3 8B and GPT 3.5 Turbo on tasks with challenging arithmetic, reaches average accuracy above that of Llama 3 8B and GPT 3.5 Turbo and close to GPT 4o + Code.
Table 2 demonstrates that arithmetic with LLMs is still challenging; state-of-the-art proprietary language models like GPT 4o achieve less than 40% accuracy on 7-digit division and fail to perform any 7-digit multiplications correctly. Open source LLMs fall farther behind, with Llama 3 8B achieving below 50% on relatively simple tasks such as 7-digit addition.
4.2 Mathematical Reasoning Problems
To test the performance of OccamLlama on longer generation tasks requiring a mix of reasoning and arithmetic capabilities, we evaluate our method and baselines on the following six benchmarks: AddSub [39], MultiArith [33], MATH401 [8], Single Eq [40], and SVAMP [41].
Because many of the arithmetic operations required in these datasets are relatively simple, we also create MultiArith Float, a modification of MultiArith in which we select problems which are arithmetically more challenging, while requiring similar levels of reasoning. To this end, we select prompts having input numbers that can be replaced with floats (e.g., $3.5$ feet or \$$\$39.95$ are reasonable but $3.5$ people is not). Furthermore, we sample input values from ranges larger than those appearing in the MultiArith dataset, in cases where it is reasonable. Float operations and larger additions and multiplications are more difficult for the baseline LLMs but do not make a difference for OccamLLM, so this dataset is particularly useful to show the advantages of this system.
Figure 3 shows the results of these evaluations. More detailed results are shown in Appendix A. On MultiArith Float and Math401, two datasets requiring challenging arithmetic, OccamLlama outperforms not only Llama 3 8B but also GPT 3.5 Turbo. On MultiArith Float, we find that OccamLlama matches the performance of GPT 4o. At the same time, most other datasets in this benchmark do not involve challenging arithmetic, meaning that Llama 3 8B is well suited to solve these tasks without assistance; most of the difficulty of these tasks lies in the reasoning rather than in the arithmetic computations. This is further supported by the fact that GPT 4o with Code Interpreter does not outperform, and in fact often underperforms, GPT 4o. As such, it is remarkable that OccamLlama can achieve comparable accuracy to Llama 3 8B even when it is trained on substantially different data and evaluated on tasks without challenging arithmetic.
In Figure 4, we show example generations from OccamLlama for both arithmetic and reasoning tasks. These generations demonstrate how the OccamLlama switch learns to balance OccamNet outputs with LLM outputs, effectively distributing the work between a reasoner (Llama) and a calculator (OccamNet). Because the language model is unaware of the OccamLlama system, its generations behave as if it possesses an interior calculator even though it is actually using a tool. In this way, we combine the benefits of a language model finetuned on arithmetic with the benefits of a language model finetuned to use code for arithmetic, all without any finetuning.
Figure 4: Examples from Llama 3 8B Instruct and OccamLlama on arithmetic and a mathematical reasoning problem from the MultiArith Float dataset. In OccamLlama, the LLM performs reasoning, the switch predicts when to use OccamNet, and OccamNet performs arithmetic operations. Inputs to OccamNet are highlighted in purple. Outputs of OccamNet are highlighted in green.
4.3 Limitations
In most of our experiments, we use a single-layer OccamNet as the symbolic network, enabling evaluation of single-operation arithmetic problems. This sometimes poses a challenge on reasoning problems, when the base language model generates compound expressions containing more than one operation, such as $3+5+7=$ or $3\cdot 20\%$ where a single-layer OccamNet cannot evaluate. We attempt to overcome this by prompting Llama to break down compound expressions into multiple steps, but we find it difficult to coerce Llama to follow these instructions. Fortunately, we observed that these compound expressions were typically simple enough for the Llama to evaluate without OccamNet. As such, in our experiments, we trained OccamLLM switch to avoid using OccamNet for compound operations. In future work, we can explore other solutions such as integrating a two-layer OccamNet as the symbolic network. Another challenge is that Llama often generates expressions in fractions or percentages, and again, we found it difficult to coerce Llama to generate only decimals; fractions and percentages also constitute compound expressions and thus limits a single-layer OccamNet’s ability to perform helpful arithmetic. We found that these issues are particularly acute in the GSM8K [42] dataset, where the expressions generated by Llama are not prevalent in the switch training data, so sometimes incorrectly triggers OccamNet, thereby degrading performance. We discuss this phenomenon in more detail in Appendix A.
[…]
5 Discussion
We presented OccamLLM, a system enabling exact and interpretable language model arithmetic in a single autoregressive step. Our method does not require modifying the weights of the underlying language model, thereby avoiding risks of catastrophic forgetting. Furthermore, our method avoids security risks arising from running code generated by a language model.
Table 3: Accuracy on multistep arithmetic.
OccamLlama Llama 3
8b Instruct
One-Step 99.9$\pm$0.1 78.1$\pm$1.3
Two-Step 98.2$\pm$0.4 57.8$\pm$1.6
Three-Step 96.1$\pm$0.6 40.2$\pm$1.6
Average 98.1$\pm$0.3 58.7$\pm$0.9
We benchmarked our method on challenging arithmetic tasks, achieving 100% accuracy whereas GPT 4o achieves only 40% performance on average. We also benchmarked our method on mathematical reasoning tasks, demonstrating that the OccamLlama switch can accurately balance the LLM for reasoning and OccamNet for arithmetic, and outperforming Llama 3 8B Instruct on tasks involving challenging arithmetic. Our work could enable smaller LLMs to be as good as much much bigger LLMs in arithmetic; moreover, integrating OccamLLM with larger LLMs like GPT 4o could further improve their arithmetic abilities without relying on a code interpreter, but we leave this for future work. Furthermore, at present, OccamLLM may not integrate with more advanced decoding techniques such as speculative decoding [38, 43]. We hope to explore these avenues in future work.
6 Broader Impact
We believe that, in addition to enabling fast, safe, and interpretable arithmetic, OccamLLM demonstrates a new paradigm for tool use. As a proof of concept for more complex tool use, we further train OccamLlama with a two layer Complete OccamNet with the primitives
$\mathcal{P}=\{\mathsf{Addition}(\cdot,\cdot),\mathsf{Subtraction}(\cdot,\cdot),\mathsf{Multiplication}(\cdot,\cdot),\mathsf{Division}(\cdot,\cdot)\},$
which enables OccamLlama to perform up to three arithmetic operations (e.g., $2\cdot 7+3/2$) in a single autoregressive step. We find that this two-layer OccamLlama can reach near 100% accuracy, even when performing three arithmetic operations in a single autoregressive step, as shown in Table 3. This demonstrates that OccamLLM can be used to perform more complex operations, including composing multiple different tools.
For future work, we plan to explore integrating other tools beyond calculators through a similar technique. This is facilitated by the fact that there are no restrictions on OccamNet’s activations; in principle, tools could be placed inside activations of OccamNet, enabling OccamNet to serve as a sort of a mixture of experts for tools. While some tools, like querying a search engine, may still be most effective when integrated into language model systems through language, we believe this work demonstrates that some tools are more effective when they can be more tightly integrated into the language model.
Acknowledgements
We would like to thank Andrew Ma and Di Luo for their thoughtful discussions.
Research was sponsored by the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
This work is also supported in part by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
Appendix
Appendix A Further Experiment Details and Results
For all training and evaluation runs, we used a single 32 GB NVIDIA Tesla V100 GPU. The LLM used for the OccamLLM system is Llama 3 8B Instruct. For each of the weight decoders and the switch described in 3.1.2 and 3.1.3, respectively, we used two-layer MLPs of input size 4096 (Llama-3-8B-Instruct hidden size), intermediate size 64 and final size equal to the number of weights in the corresponding OccamNet layer or switch.
A.1 Training Dataset
We created a training dataset consisting of a 80,000 examples split in 40,000 single queries and 40,000 sequences of concatenated queries. In the first case, a single prompt of those described in 3.2 was sampled and formatted in the Llama 3 8B Instruct chat template, and the result of the expression was added as the output of the model. 40% of the sampled prompts correspond to simple arithmetic, concretely +,-,*,/. We sampled from various input value ranges. We sampled integers and floating point numbers uniformly from $(-20000,20000)$ and $(-1000,1000)$, integers only from $(-100,100)$ and $(-10,10)$ and floating points only from $(-1,1)$.
Another 40% corresponds to complex arithmetic involving square roots, logarithms, exponentials, trigonometric functions and computing the power of one number to another one. For the square root and the logarithm, we sampled integers and floats uniformly from $(0.01,20000)$ and $(0.01,100)$. For the exponential, we sampled integers and floats from $(-10,10)$. For the powers, both base an exponents were sampled as integers or floats from $(1,25)$ and $(-6,6)$, respectively.
The remaining 20% corresponds to reasoning prompts, either single or multi step problems. The inputs were sampled with various ranges, sometimes as floats and sometimes as integers, depending on the context of the problem. To train this model, the multiple-query examples never finished in a multi reasoning problem, as these would require more than one layer of OccamNet depth to be predicted accurately.
We first iterated the 80,000 examples prepending "Answer = " to the assistant response, thus training OccamNet to predict the result after the "= ". Next, we validated the model on out-of-distribution examples, were it was forced to predict the result right at the beginning of the assistant response. We noticed that the accuracy on this task was improving during training, but after the full dataset was iterated it still didn’t perform as well as when evaluated in-distribution. Therefore, we continued to train the model using examples of the same dataset but with no appendix to the assistant response, and it rapidly learnt the new task. We stopped at 28,000 iterations in this second stage.
To train the switch, we create a dataset of 50,000 examples. Each example consists of a series of variable length input-output pairs sampled from a collection of prompts, which are formatted in the Llama 3 8B Instruct chat template in a conversational style. For each pair, the tokens previous to the numbers that should be computed using OccamNet, which are the ones that the switch should not route to the LLM, are marked with a 1, and all the rest are marked with a 0.
25% of these pairs are created by taking one of the arithmetic expressions described above as input, and using as output either “Answer = ”, the same expression repeated or nothing at all.
70% of the pairs come from a collection of manually created and labeled examples, which illustrate in which cases the switch should route to OccamNet and, importantly, in which cases it shouldn’t. This collection was deisgned to cover a wide variety of situations where the LLM might need to use OccamNet for computations. Furthermore, it includes cases where the LLM should avoid calling OccamNet because doing so would produce a bad prediction. This is the case, for example, of instances where the LLM attempts to add three numbers simultaneously. If it were to use the 1-layer OccamNet, which can take 2 inputs at most, the result would be incorrect.
The remaining 5% of the prompts come from multi-step reasoning problems. The output for these is not a full response, but only “The answer is ”.
For the two-layer OccamNet run, we generated a large set of programatically generated prompts of the form 3+974=3+97$\cdot$-4=3 + 97 ⋅ - 4 =, with the Llama 3 8B chat template applied.
A.2 Training Hyperparameters
For all 1-layer OccamNet training runs, we used a batch size of $1$, a learning rate of $46e-$ 4 and a weight decay parameter of $0.01$. We use the gradient accumulation technique to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We use 1000 samples from OccamNet per token.
For the 2-layer OccamNet run, we use a learning $41e-$ 4, we used a batch size of $1$, a learning rate of $46.e-$ 4 and a weight decay parameter of $0.01$. We use the gradient accumulation technique to achieve an effective batch size of $8$. We used a constant learning rate scheduler. We use 50,000 samples from OccamNet per token.
A.3 Prompting
For for the division arithmetic tasks, we found that the language models often did not return decimals. As such, we appended “Give the answer in decimals.” to these prompts. Similarly, for the trigonometric functions evaluations, we explicitly ask the language models to take the input as radians, by formatting the prompts as "cos(X rad) =".
For some models, we provide system prompting to guide the model toward the correct behavior. We break down prompting by model below:
Llama 2/3:
We did not provide a system prompt for the arithmetic tasks. For the reasoning tasks, we used the system prompt “Solve step by step.”
GPT 3.5 Turbo:
We do not use a system prompt for GPT 3.5 Turbo.
GPT 4o:
We did not use a system prompt, except for on the MATH401 dataset, where we noticed that GPT 4o was returning fractions instead of decimals. As such, on MATH401 we used the system prompt “Give your answer in decimals.”
GPT 4o + Code:
We used the system prompt “Write and run code to answer math questions. Do not format numbers. Give all answers in decimals.”
OccamLlama:
We experimented with OccamLlama prompts
A.4 Generation parameters
For OccamLlama, Llama 2 7B and Llama 3 8B, we use the default values of $T=0.6$ and Top-P $=0.9$. For GPT 3.5 Turbo, GPT 4o, and GPT 4o with Code Interpreter, we use the default values of $T=1.0$ and Top-P $=1.0$.
A.5 Experimental Results
Tables 4 and 5 show in more detail the accuracy of OccamLlama and other baselines on arithmetic and reasoning tasks. We measure accuracy as described in the main text.
We note here that on reasoning tasks with challenging arithmetic, such as Multiarith Float and MATH401, OccamLlama performs on-par with or even outperforms GPT 4o and GPT 4o Code. In fact, on MATH401, OccamLlama is nearly a 10 percentage points more accurate than Llama.
The only dataset on which OccamLlama struggles is GSM8K [42]. We beleive the reason for this is primarily because this data involves a large number of fractions and percentages, which Llama does not convert to decimal format. As such, Llama often calls OccamNet with expressions such as “multiplying by 3/4 gives” or “5% of this gives.” Because the switch is not trained on many examples like these in which the number is not in decimal format, it does not realize that OccamNet cannot be used in these cases. So, the switch triggers OccamNet, which is capable of performing the correct operation because it requires two steps. Future work could address this issue by training the router with more data on this type of situation or by training an OccamLlama with a two layer OccamNet.
Relative error is another important metric that complements accuracy. It measures by how much did the answer miss the true result. For two models with a similar accuracy metric, the relative error they achieve can be very different. Table 6 shows the relative error for the arithmetic experiments.
Table 4: Accuracy on arithmetic tasks. Higher is better. Bold indicates best performance.
OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
7B Chat 8b Instruct Turbo Code
Addition (3) 100.0$\pm$0.0 70.9$\pm$1.4 97.1$\pm$0.5 98.8$\pm$0.3 100.0$\pm$0.0
Addition (5) 100.0$\pm$0.0 55.9$\pm$1.6 77.1$\pm$1.3 92.5$\pm$0.8 99.2$\pm$0.3
Addition (7) 100.0$\pm$0.0 19.2$\pm$1.2 44.9$\pm$1.6 65.2$\pm$1.5 95.7$\pm$0.6 100.0$\pm$0.0
Subtraction (3) 100.0$\pm$0.0 49.7$\pm$1.6 95.2$\pm$0.7 94.0$\pm$0.8 98.7$\pm$0.4
Subtraction (5) 100.0$\pm$0.0 22.9$\pm$1.3 58.8$\pm$1.6 86.3$\pm$1.1 92.6$\pm$0.8
Subtraction (7) 100.0$\pm$0.0 8.7$\pm$0.9 34.4$\pm$1.5 59.8$\pm$1.6 85.6$\pm$1.1 99.5$\pm$0.5
Multiplication (3) 100.0$\pm$0.0 4.6$\pm$0.7 16.8$\pm$1.2 49.2$\pm$1.6 76.9$\pm$1.3
Multiplication (5) 100.0$\pm$0.0 0.0$\pm$0.0 0.1$\pm$0.1 0.4$\pm$0.2 4.6$\pm$0.7
Multiplication (7) 100.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 99.0$\pm$0.7
Division (3) 100.0$\pm$0.0 20.8$\pm$1.3 71.7$\pm$1.4 50.5$\pm$1.6 78.2$\pm$1.3
Division (5) 100.0$\pm$0.0 7.4$\pm$0.8 48.1$\pm$1.6 15.7$\pm$1.2 51.0$\pm$1.6
Division (7) 100.0$\pm$0.0 2.8$\pm$0.5 35.3$\pm$1.5 10.7$\pm$1.0 38.6$\pm$1.5 100.0$\pm$0.0
Square Root (3) 100.0$\pm$0.0 1.2$\pm$0.3 14.8$\pm$1.1 47.1$\pm$1.6 69.3$\pm$1.5
Square Root (5) 100.0$\pm$0.0 0.2$\pm$0.1 1.3$\pm$0.4 11.9$\pm$1.0 23.6$\pm$1.3
Square Root (7) 100.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.9$\pm$0.3 18.6$\pm$1.2 100.0$\pm$0.0
Exponential 100.0$\pm$0.0 0.3$\pm$0.2 3.1$\pm$0.5 12.5$\pm$1.0 23.2$\pm$1.3 100.0$\pm$0.0
Logarithm 100.0$\pm$0.0 0.1$\pm$0.1 0.0$\pm$0.0 17.1$\pm$1.2 21.3$\pm$1.3 100.0$\pm$0.0
Sine 100.0$\pm$0.0 7.6$\pm$0.8 7.0$\pm$0.8 13.4$\pm$1.1 39.3$\pm$1.5 100.0$\pm$0.0
Cosine 100.0$\pm$0.0 0.8$\pm$0.3 1.5$\pm$0.4 6.7$\pm$0.8 32.8$\pm$1.5 100.0$\pm$0.0
Average 100.0$\pm$0.0 14.4$\pm$0.3 32.0$\pm$0.3 38.6$\pm$0.4 55.2$\pm$0.4 99.8$\pm$0.1
Table 5: Accuracy on reasoning tasks. Higher is better. Bold indicates best non-GPT-4o performance.
OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
7B Chat 8b Instruct Turbo Code
AddSub 91.4$\pm$1.4 78.0$\pm$2.1 93.4$\pm$1.2 95.4$\pm$1.1 97.0$\pm$0.9 97.5$\pm$1.1
GSM8K 68.8$\pm$1.3 36.0$\pm$1.3 79.8$\pm$1.1 84.8$\pm$1.0 96.1$\pm$0.5 94.0$\pm$1.7
MultiArith 99.5$\pm$0.3 76.0$\pm$1.7 99.8$\pm$0.2 97.2$\pm$0.7 99.7$\pm$0.2 99.5$\pm$0.5
MultiArith Float 96.0$\pm$0.8 23.3$\pm$1.7 57.3$\pm$2.0 77.3$\pm$1.7 96.2$\pm$0.8 89.5$\pm$2.2
MATH401 86.3$\pm$1.7 43.9$\pm$2.5 60.3$\pm$2.4 63.1$\pm$2.4 76.6$\pm$2.1 78.0$\pm$2.9
Single Eq 91.3$\pm$1.2 79.1$\pm$1.8 96.3$\pm$0.8 97.8$\pm$0.6 98.0$\pm$0.6 99.0$\pm$0.7
SVAMP 87.9$\pm$1.0 61.5$\pm$1.5 86.3$\pm$1.1 87.8$\pm$1.0 96.2$\pm$0.6 96.5$\pm$1.3
Average 85.7$\pm$0.5 53.3$\pm$0.7 82.1$\pm$0.6 86.5$\pm$0.5 95.2$\pm$0.3 93.4$\pm$0.7
Table 6: Relative error on arithmetic tasks. Higher is Better. Bold indicates best performance.
OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
7B Chat 8b Instruct Turbo Code
Addition (3) 0.0$\pm$0.0 50.5$\pm$10.9 3.2$\pm$1.9 0.3$\pm$0.1 0.0$\pm$0.0
Addition (5) 0.0$\pm$0.0 113.0$\pm$21.1 23.7$\pm$4.0 4.6$\pm$1.8 0.0$\pm$0.0
Addition (7) 0.0$\pm$0.0 310.3$\pm$97.0 78.1$\pm$16.2 4.0$\pm$1.4 1.0$\pm$0.9 0.0$\pm$0.0
Subtraction (3) 0.0$\pm$0.0 66.2$\pm$18.7 4.1$\pm$0.8 3.8$\pm$0.7 0.4$\pm$0.1
Subtraction (5) 0.0$\pm$0.0 173.6$\pm$67.1 29.4$\pm$4.3 38.3$\pm$16.5 3.5$\pm$0.6
Subtraction (7) 0.0$\pm$0.0 222.3$\pm$54.4 65.6$\pm$12.9 44.6$\pm$31.3 5.4$\pm$0.7 0.3$\pm$0.3
Multiplication (3) 0.0$\pm$0.0 7.6$\pm$0.7 1.9$\pm$0.5 1.8$\pm$0.4 0.1$\pm$0.1
Multiplication (5) 0.0$\pm$0.0 84.9$\pm$0.9 46.7$\pm$1.7 19.2$\pm$3.7 1.8$\pm$0.4
Multiplication (7) 0.0$\pm$0.0 98.9$\pm$0.2 74.9$\pm$1.8 90.1$\pm$24.2 4.4$\pm$0.6 1.0$\pm$0.7
Division (3) 0.1$\pm$0.0 1346.2$\pm$275.4 1.3$\pm$0.9 1.1$\pm$0.3 0.0$\pm$0.0
Division (5) 0.2$\pm$0.1 174156.6$\pm$31687.6 9.5$\pm$1.8 0.7$\pm$0.2 0.1$\pm$0.0
Division (7) 0.1$\pm$0.0 22032920.9$\pm$3642549.7 225.3$\pm$142.1 0.3$\pm$0.1 0.0$\pm$0.0 0.0$\pm$0.0
Square Root (3) 0.0$\pm$0.0 8.9$\pm$1.1 1.1$\pm$0.3 0.2$\pm$0.0 0.0$\pm$0.0
Square Root (5) 0.0$\pm$0.0 72.8$\pm$4.8 12.6$\pm$1.7 0.1$\pm$0.0 0.0$\pm$0.0
Square Root (7) 0.0$\pm$0.0 207.8$\pm$21.8 15.4$\pm$1.4 8.8$\pm$0.9 4.8$\pm$2.0 0.0$\pm$0.0
Exponential 0.3$\pm$0.0 422.6$\pm$82.6 11.7$\pm$0.7 2.3$\pm$0.9 0.1$\pm$0.0 0.0$\pm$0.0
Logarithm 0.0$\pm$0.0 138.3$\pm$11.9 40.2$\pm$1.3 6.4$\pm$4.0 0.1$\pm$0.0 0.0$\pm$0.0
Sine 0.1$\pm$0.0 515.5$\pm$68.1 279.2$\pm$21.9 125.4$\pm$12.3 3.2$\pm$0.7 0.1$\pm$0.0
Cosine 0.1$\pm$0.0 581.9$\pm$199.7 346.1$\pm$80.4 188.5$\pm$22.5 2.8$\pm$0.5 0.0$\pm$0.0
Average 0.1$\pm$0.0 1169210.8$\pm$194953.4 66.9$\pm$8.8 28.4$\pm$2.7 1.5$\pm$0.1 0.2$\pm$0.1
Appendix B Alternative Architectures and Losses
B.1 Alternative Architectures
As discussed in the main text, although OccamLLM works most naturally with OccamNet, it can also work with other symbolic architectures such as the EQL network [44, 45], or architectures that can represent probability distributions over symbolic expressions, such as transformers [31] or recurrent neural networks (RNNs) [32].
However, in practice we believe OccamNet is the most effective architecture for this use case. We find that because EQL does not represent a probability distribution over functions, it easily gets stuck in local minima.
Regarding transformers and RNNs, we believe that OccamNet possesses a key advantage of being interpretable; by looking at the weights, it is possible for a human to see what the architecture is doing. We beleive that this interpretability will also make it easier for a decoder to initialize properly, as opposed to an RNN or transformer where the relation between the weights and the probability distribution is substantially more complex. This leads us to our key point. Transformers and RNNs are effective for modeling complex multimodal distributions, but for this problem, we want to select a single function for each token, so the extra expressivity of these models is unneeded and likely detrimental to performance. We beleive that OccamNet, being a much simpler architecture, enables better parameter efficiency and performance.
B.2 Alternative Losses
In this section we discuss alternative possible losses and how we arrived at the loss in Equation 3.
We considered two loss functions which are natural when optimizing a probability distribution: 1) a cross entropy loss, and 2) a reinforcement learning loss.
Cross entropy losses are effective, but they require access to the ground-truth distribution. As such, this approach may not scale well for OccamNet.
Reinforcement learning losses, on the other hand are effective for exploring large search spaces. We find that using REINFORCE with a delta function for our reward is most effective because it mimics a cross entropy loss. We rescale the loss by the total reward for all the samples to ensure that correct functions sampled only a few times still receive a large training step update.
Appendix C Background on OccamNet
Much of this section is modified from [11].
We divide this section into the following subsections:
  1. 1.
    In Section C.1, we describe OccamNet’s architecture in more detail.
  2. 2.
    In Section C.2, we describe OccamNet’s sampling process.
  3. 3.
    In Section C.3, we describe OccamNet’s probability distribution.
  4. 4.
    In Section C.4, we describe OccamNet’s initialization process.
C.1 OccamNet Architecture
As described in the main text, we start from a predefined collection of $N$ primitive functions $\mathcal{P}$. OccamNet represents a distribution over compositions of functions in $\mathcal{P}.$ From now on, we denote the $i$th primitive function in the $l$th layer as $\phi_{i}^{(l)}$. We 0 index the primitives but 1 index the layers, because we treat the inputs as the 0th layer. So, for example, in Figure 2a, $\phi_{2}^{(1)}=\phi_{2}^{(2)}=\phi_{2}^{(3)}=\mathsf{sin}$.
Figure 5: A demonstration of the dropped connections from sampled paths in OccamNet. All light grey paths are dropped from the final symbolic form of the sampled function because they are not directly connected to the outputs. These paths are unnecessarily computed during OccamNet’s training process, leading to potential slowdowns in training.
Each OccamNet layer consists of two sublayers, which we denote the arguments and image sublayers, shown in Figure 5a. For an $L$-layer OccamNet, each of these sublayers is reproduced $L$ times. The $l$th softmax layer connects the $(l-1)$th image layer with the $l$th arguments layer. For $1\leq l\leq L$, we denote the $l$th arguments sublayer hidden state as $\widetilde{\mathbf{h}}^{(l)}$ and the $l$th image sublayer hidden state as $\mathbf{h}^{(l)}$. So, $\widetilde{\mathbf{h}}^{(2)}$ would represent the We further write
$\mathbf{\widetilde{h}}^{(l)}=\left[\widetilde{h}^{(l)}_{1},\dots,\widetilde{h}^{(l)}_{M^{(l)}}\right],\;\;\mathbf{h}^{(l)}=\left[h^{(l)}_{1},\dots,h^{(l)}_{N^{(l)}}\right],$ (4)
where
$M^{(l)}=\sum_{0\leq k\leq N^{(l)}}\alpha\left[\phi_{k}^{(l)}\right]$
and $\alpha[\phi]$ is the arity of function $\phi.$ We also define $\mathbf{h}^{(0)}$ to be the input layer (an image sublayer) and $\mathbf{\widetilde{h}}^{(L+1)}$ to be the output layer (an arguments sublayer).
In a standard OccamNet layer, each primitive is repeated exactly once in each layer. However, in Complete OccamNet, each primitive in the $l$th layer is repeated $A^{L-l}$ times, where $A$ is the maximum arity of the primitives. This is shown in Figure 6 in the transition from 6a) to 6b). Complete OccamNet also concatenates each image layer to the next image layer, as shown in Figure 6c.
C.2 Sampling from OccamNet
In this section, we more carefully describe OccamNet’s sampling process. For each softmax layer, we sample a connection to each arguments layer node from the softmax of the weights. In particular, if $\mathbf{w}_{i}^{(l)}$ are the weights of the $l$th softmax layer leading to the $i$th node of the $l$th argument’s layer, when we sample we produce a sparse matrix
$\mathsf{SAMPLE}\left(\begin{bmatrix}{\mathsf{softmax}(\mathbf{w}_{1}^{(l)})}\\ \vdots\\ {\mathsf{softmax}(\mathbf{w}_{M^{(l)}}^{(l)})}\end{bmatrix}\right)$ (5)
where the $\mathsf{SAMPLE}$ function samples a one-hot row vector for each row based on the categorical probability distribution defined by $\mathsf{softmax}(\mathbf{w})$. To evaluate a forward pass through this sample, we simply evaluate a forward pass through the network, treating the sampled sparse matrices from the softmax layers as the weights of linear layers:
$\widetilde{\mathbf{h}}^{(l)}=\begin{bmatrix}\widetilde{h}_{1}^{(l)}\\ \vdots\\ \widetilde{h}_{M^{(l)}}^{(l)}\end{bmatrix}\equiv\mathsf{SAMPLE}\left(\begin{bmatrix}{\mathsf{softmax}(\mathbf{w}_{1}^{(l)})}\\ \vdots\\ {\mathsf{softmax}(\mathbf{w}_{M^{(l)}}^{(l)})}\end{bmatrix}\right)\mathbf{h}^{(l-1)},$ (6)
To complete the picture of the forward pass, we formalize how we deal with activations accepting multiple inputs. We define the action of the activation functions as follows:
$h^{(l)}_{i}=\phi_{i}^{(l)}\left(\widetilde{h}_{j}^{(l)},\dots,\widetilde{h}_{j+\alpha(\phi_{i})-1}^{(l)}\right),\;\;j=\sum_{0\leq k<i}\alpha(\phi_{k}^{(l)}).$ (7)
Figure 6: Skip connections. Nodes are color coded with lines indicating the origin of the reused neurons.
C.3 OccamNet’s Probability Distribution
OccamNet parametrizes a probability distribution over all functions which it can sample. In particular, when OccamNet samples a function, it is really sampling a directed acyclic graph (DAG) which defines a computational path to compute a function. The probability of sampling a function is equal to the product of the probabilities of the connections in the DAG which are connected to the output node.
When using OccamLlama for inference, we select the maximum probability function by sampling 100 functions from OccamNet, evaluating their probabilities as described above and selecting the maximum one.
C.4 Initialization
This section describes how we calculate $\mathbf{W}_{i}^{(0)}$ from the main text.
The algorithm to initialize all functions with equal probability establishes as an invariant that, after assigning the weights up to the $l$th layer, all paths leading to a given node in the $l$th argument layer have equal probabilities. Then, each argument layer node has a unique corresponding probability, the probability of all paths up to that node. We denote the probability of the $i$th node in the $l$th argument sublayer as $\widetilde{p}_{i}^{(l)},$ because it is the probability of any path leading to the $i$th node in the $l$th argument sublayer. Because each argument layer node has a corresponding probability, each image layer node must also have a unique corresponding probability, which, for the $i$th node in the $l$th image sublayer, we denote as $p_{i}^{(l)}.$ Again, we use the notation $p_{i}^{(l)}$ because this is the probability of any path leading to the $i$th node in the $l$th image sublayer. These image layer probabilities are given by
[…]
Our algorithm starts with input layer, or the 0th image layer. Paths to any node in the input layer have no edges so they all have probability 1. Thus, we initialize $p_{i}^{(0)}=1$ for all $i$. As the algorithm iterates through all subsequent $T$-Softmax layers, the invariant established above provides a system of linear equations involving the desired connection probabilities, which the algorithm solves. The algorithm groups the previous image layer according to the node probabilities, obtaining a set of ordered pairs $\{({p^{\prime}}_{a}^{(l)},n_{a}^{(l)})\}_{i=a}^{k}$ representing $n_{a}^{(l)}$ nodes with probability ${p^{\prime}}_{a}^{(l)}$ in the $l$th layer. Note that if two image nodes have the same probability $p_{i}^{(l)}=p_{j}^{(l)}$, then the edges between any argument node in the next layer and the two image nodes must have the same probability in order to satisfy the algorithm’s invariant: $p_{i}^{(l,k)}=p_{j}^{(l,k)}$. Then, we define ${p^{\prime}}_{a}^{(l,i)}$ as the probability of the edges between the image nodes with probability ${p^{\prime}}_{a}^{(l)}$ and the $i$th argument $P$-node of the $l$th layer. The probabilities of the edges to a given $P$-node sum to 1, so for each $j$, we must have $\sum_{a}n_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)}=1.$ Further, the algorithm requires that the probability of a path to a $P$-node through a given connection is the same as the probability of a path to that $P$-node through any other connection. The probability of a path to the $i$th $P$-node through a connection with probability ${p^{\prime}}_{a}^{(l,i)}$ is ${p^{\prime}}_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)},$ so we obtain the equations ${p^{\prime}}_{0}^{(l)}{p^{\prime}}_{0}^{(l,i)}={p^{\prime}}_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)},$ for all $a$ and $i$. These two constraints give the vector equation
$\begin{bmatrix}n_{0}^{(l)}&n_{1}^{(l)}&n_{2}^{(l)}&\cdots&n_{k}^{(l)}\\ {p^{\prime}}_{0}^{(l)}&-{p^{\prime}}_{1}^{(l)}&0&\cdots&0\\ {p^{\prime}}_{0}^{(l)}&0&-{p^{\prime}}_{2}^{(l)}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {p^{\prime}}_{0}^{(l)}&0&0&\cdots&-{p^{\prime}}_{k}^{(l)}\end{bmatrix}\begin{bmatrix}{p^{\prime}}_{0}^{(l,j)}\\ {p^{\prime}}_{1}^{(l,j)}\\ {p^{\prime}}_{2}^{(l,j)}\\ \vdots\\ {p^{\prime}}_{k}^{(l,j)}\end{bmatrix}=\begin{bmatrix}1\\ 0\\ 0\\ \vdots\\ 0\end{bmatrix},$ (9)
for all $1\leq j\leq M$. The algorithm then solves for each ${p^{\prime}}_{a}^{(l,j)}.$
After determining the desired probability of each connection of the $l$th layer, the algorithm computes the SPL weights ${\mathbf{w}^{\prime}}^{(l,j)}$ that produce the probabilities ${p^{\prime}}_{a}^{(l,j)}$. Since there are infinitely many possible weights that produce the correct probabilities, the algorithm sets ${w^{\prime}}_{0}^{(l,j)}=0.$ Then, the algorithm uses the softmax definition of the edge probabilities to determine the required value of $\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})$:
${p^{\prime}}_{0}^{(l,j)}$ $=\frac{\exp({w^{\prime}}_{0}^{(l,j)}/T^{(l)})}{\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})}$
$=\frac{1}{\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})}$
so
$\sum_{a=1}^{k}\exp({w^{\prime}}_{a}^{(l,j)}/T)=1/{p^{\prime}}_{0}^{(l,j)}.$
Substituting this equation into the expression for the other probabilities gives
${p^{\prime}}_{a}^{(l,j)}$ $=\exp({w^{\prime}}_{a}^{(l,j)}/T^{(l)})/\left(\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})\right)$
$={p^{\prime}}_{0}^{(l,j)}\exp({w^{\prime}}_{a}^{(l,j)}/T^{(l)}).$
Solving for ${w^{\prime}}_{i}^{(l,j)}$ gives
${w^{\prime}}_{a}^{(l,j)}=T^{(l)}\log\left({p^{\prime}}_{a}^{(l,j)}/{p^{\prime}}_{0}^{(l,j)}\right),$ (10)
which the algorithm uses to compute ${w^{\prime}}_{i}^{(l,j)}$.
After determining the weights ${w^{\prime}}_{a}^{(l,j)}$ the algorithm assigns them to the corresponding $w_{a}^{(l,j)}.$ In particular, if the $i$th image node has probability ${p^{\prime}}_{a}^{(l)},$ the weights of edges to the $i$th node are given by $w_{i}^{(l,j)}={w^{\prime}}_{k}^{(l,j)},$ for all $j.$ The algorithm then determines the values of $\widetilde{p}_{j}^{(l+1)},$ given by $\widetilde{p}_{j}^{(l+1)}=p_{i}^{(l)}p_{i}^{(l,j)}.$ Finally, the algorithm determines $p_{i}^{(l+1)}$ using Equation 8 and repeats the above process for subsequent layers until it reaches the end of the network.
In summary, the algorithm involves the following steps:
  1. 1.
    Set $l=0$.
  2. 2.
    Set $p_{i}^{(l)}=1$.
  3. 3.
    Increment $l$ by 1.
  4. 4.
    Compute $\{({p^{\prime}}_{a}^{(l)},n_{a}^{(l)})\}_{i=a}^{k}$ and use Equation 9 to compute ${p^{\prime}}_{a}^{(l,j)}.$
  5. 5.
    Set ${w^{\prime}}_{a}^{(l,j)}$ according to Equation 10.
  6. 6.
    If $l<L+1$, Compute $\widetilde{p}_{i}^{(l+1)}$ and $p_{i}^{(l+1)}$.
  7. 7.
    Return to step 3 until $l=L+1.$
This algorithm efficiently equalizes the probabilities of all functions in the network. In practice, however, we find that perfect equalization of functions causes activation functions with two inputs to be highly explored. This is because in this section we have implicitly been using the approximate probability $\tilde{q}.$ This probability underestimates many functions that include activation functions with two or more inputs because these functions are those which can use a node multiple times in their DAG. As a result, although all functions will have an equal $\tilde{q}$, some functions with multiple inputs will have larger $q$ than other functions, and $q$ is what determines the probability of being sampled. However, we find that this initialization is still effective for approximately equalizing the probabilities at the start of training.